OpenStudy (anonymous):

find the area of a triangle with the height of 8 and a base of x-5

OpenStudy (campbell_st):

the formula for the area of a triangle is \[A = \frac{1}{2} \times base \times height\] substitute your known values and evaluate

OpenStudy (anonymous):

so it's 4x-20 but do you need ^2 at the end?

OpenStudy (campbell_st):

you need the units of measurement expressed as square units e.g. cm^2 or ft^2